significance for bone density scores are normally distributed with a mean of 0 and a standard deviation of I, find the percentage of score that are a. significant high(or at least 2 standard deviations above the mean
Answer #1

P(X > 2)

= P((X - )/> (2 - )/)

= P(Z > (2 - 0)/1)

= P(Z > 2)

= 1 - P(Z < 2)

= 1 - 0.9772

= 0.0228 = 2.28%